As Tony the fisherman reels in a "big one," he turns the spool on his fishing reel at the rate of 3.0 complete revolutions every second.If the radius of the reel is 2.5 cm, what is the linear speed of the fishing line as it is reeled in?How would your answer to part A change if the radius of the reel were doubled?
Answer #1

2*Pi*r *3 = speed/second

Put values..
2*3.142*2.5 *3 = 47.13 cm/s

if doubled r = 5
2*3.142*5*3 = 94.26 cm/s

It has doubled.

I hope help you !!